Anthony leaves Kingstown and drives 160 miles to Albany. He leaves at 2:00 p.m. At 2:15p.m., Emily leaves Albany and drives to Kingstown at 40 m.p.h. If Anthony is driving 45 m.p.h., at what time do they pass each other on the road?
Answer by jorel1380(3719)

2:15-2:00=1/4 hour
1/4(45)=11.25 miles
160-11.25=148.75 miles
148.75/(40+45)=148.75/85=1.75 hours
They pass each other at 1-3/4 hours after 2:15, or at 4:00 p.m...